# VRSBench: A Versatile Vision-Language Benchmark Dataset for Remote Sensing Image Understanding

### v1

@@ 1 Introduction
Remote sensing models seek to understand the Earth’s surface using imagery captured from overhead, offering a unique perspective of our physical world. This technique is instrumental in various applications, such as land use mapping, urban planning, precision agriculture, disaster management, etc. In the past few years, the success of large vision-language models (LVMs)[1, 2, 3, 4, 5] in natural scenes has inspired a trend of applying LVMs to remote sensing [6]. Recent efforts have explored LVMs for various remote sensing image understanding tasks, including scene classification [7, 8], image captioning [9, 10, 11, 12, 13, 14, 15, 16, 17], visual grounding [18, 19, 20], visual question answering (VQA) [21, 22, 23, 24, 25, 26], and general-purpose models [27, 28], etc.
However, directly applying LVMs to remote sensing images presents challenges. LVMs are typically trained on internet data, which differs significantly from remote sensing data. Remote sensing images often feature very small objects (sometimes only 10 pixels) and require complex spatial reasoning from an overhead view. Building effective LVMs for remote sensing requires large-scale, high-quality datasets tailored to this field. Recent works[27, 28, 29] have attempted to train LVMs with a combination of existing text-enriched remote sensing data, achieving reasonable performance. However, further improvements are limited by the current vision-language datasets in remote sensing, which have the following limitations:
@@
The key contributions of our work are summarized as follow:
@@
  • VRSBench provides large-scale human-verified annotations that feature several advantages: 1) it provides a large-scale collection of human-verified, high-quality captions rich in object details; 2) it offers more realistic object refers in which each referring sentence unambiguously identifies an object among multiple similar ones within the same category; 3) it features a diverse collection of open-set question-answer pairs in natural language.
  • We develop three benchmarks based on our VRSBench dataset, including detailed image caption, visual grounding, and visual question answering, and evaluate the performance of several state-of-the-art LVMs.
@@
#Captions Details #Refers OBB #VQAs Open-set
@@
Table 1: Comparison between existing remote sensing vision-language datasets and our VRSBench dataset. Values in parentheses in the Caption column indicate the average word length of captions. OBB denotes orientated bounding box. A small portion of question-answer pairs in RSIVQA are annotated by human annotators.
@@ 2.1 Attribute Extraction
In this study, we utilize two prominent open-access object detection datasets, DOTA-v2 [33] and DIOR [34], to develop our VRSBench dataset. Due to the unavailability of test labels for DOTA-v2, we incorporate only its training and validation sets. We divide each image into patches measuring 512 × 512 pixels. Notably, each image patch from DOTA-v2 contains, on average, 14.2 instances, while each patch from the DIOR dataset averages only 3.3 instances. This higher instance density in DOTA-v2 offers a more challenging and diverse training environment compared to existing remote sensing visual grounding datasets, such as DIOR-RSVG [19] and RRSIS-D [35], that are typically sourced from the DIOR dataset. Moreover, the visual grounding task predominantly involves identifying horizontal bounding boxes (HBB) based on referential descriptions. By constructing our dataset upon the framework of DOTA-v2, VRSBench facilitates the grounding of objects with orientated bounding boxes (OBB), thereby extending the capabilities of traditional visual grounding mothods.
@@ 4.1 Benchmark Overview
For the above three tasks, we benchmark state-of-the-art models, including LLaVA-1.5 [37], MiniGPT-v2 [38], and GeoChat [28], to demonstrate the potential of LVMs for remote sensing image understanding. LLaVA-1.5 [37], MiniGPT-v2 [38], Mini-Gemini [2], and GeoChat [28] are generalist models that are naturally designed for general-proposed image understanding. We, therefore, report the performance of these methods under joint training of all three tasks, i.e., image captioning, visual grounding, and VQA.
Note that all these comparison methods include a multi-stage training process. To ensure a fair comparison, we reload the models that are initially trained on large-scale image-text alignment datasets, and then finetune each method using the training set of our VRSBench dataset. We employ CLIP-ViT(L-14) [39] as the vision encoder and use the Vicuna-7B model [40] as the LLM. For LLaVA-1.5 [37], Mini-Gemini [2], and GeoChat [28], we adhere to the original model specifications, utilizing two MLP layers with GeLU activation [41]. For MiniGPT-v2[38], we implement a single MLP layer as described in the original paper. For each comparing method, we finetune the model on the training set of our VRSBench dataset for 5 epochs. Following GeoChat [28], we use LoRA [42] finetuning to finetune all comparing methods, with a rank of 64. For Mini-Gemini [2], we use an input resolution of 336 for the vision tower and 768 for the auxiliary vision tower. To understand the benefit of fientuning on VRSBench, we include the baseline GeoChat [28] without training on our VRSBench dataset for comparison.
@@ 4.2 Detailed Image Caption
Evaluation metrics. For model evaluation, we follow standard practices by utilizing a set of established metrics including BLEU [43], ROUGE_L [44], METEOR [45], and CIDEr [46]. For BLEU, we consider n-gram precision with n values of 1, 2, 3, and 4. We also report average caption lengths to assess the details of generated captions.
Results. Table 3 shows the comparative performance of different methods in detailed image captioning of our VRSBench dataset. As demonstrated in the table, the baseline GeoChat model, when not finetuned on the VRSBench dataset, exhibits significantly poorer performance compared to models that have been finetuned on VRSBench. The LLaVA-1.5 [37] model that has undergone fine-tuning on VRSBench achieves the highest performance, reaching a BLEU-1 score of 48.1 and a CIDEr score of 33.9. Moreover, the generated captions have an average word length of 52, which closely approximates the average length of ground truth captions. Qualitative results can be found in the supplementary. It should be noted that detailed image caption is a more challenging task than conventional image caption, therefore, the performance falls far behind. More advanced vision-language modeling techniques are desired to handle this challenging task.
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE_L CIDEr Avg_L
GeoChat w/o ft [28] 13.9 6.6 3.0 1.4 7.8 13.2 0.4 36
GPT-4V [32] 37.2 22.5 13.7 8.6 20.9 30.1 19.1 67
MiniGPT-v2 [38] 36.8 22.4 13.9 8.7 17.1 30.8 21.4 37
LLaVA-1.5 [37] 48.1 31.5 21.2 14.7 21.9 36.9 33.9 49
GeoChat [28] 46.7 30.2 20.1 13.8 21.1 35.2 28.2 52
Mini-Gemini [47] 47.6 31.1 20.9 14.3 21.5 36.8 33.5 47
Table 3: Detailed image caption performance on VRSBench dataset. Avg_L denotes the average word length of generated captions. Boldface and underline indicate the best and second-best performance.
@@ 4.3 Visual Grounding
Results. Table 3 shows the visual grounding performance of different methods on our VRSBench dataset. From the table, the model finetuned on the VRSBench training set significantly outperforms the baseline GeoChat model without finetuning. Furthermore, all models demonstrate superior performance in tasks involving unique object referring compared to non-unique object referring. This superiority is expected, as it is generally easier to localize objects uniquely identified within their categories than to differentiate among multiple instances within the same category. Note that even though MiniGPT-v2 gets worse overall grounding performance, it performs better at grounding non-unique objects.
@@
More importantly, even the best-performing GeoChat model fails to achieve satisfactory performance levels, with a grounding accuracy of 39.6% at a threshold of 0.5. This shortfall is attributed to the demanding scenarios presented in the VRSBench dataset, which includes multiple instances of the same category as the target object. This highlights the necessity for more advanced vision grounding techniques to effectively tackle these complexities. Qualitative results are provided in the supplementary.
@@
LLaVA-1.5 [37] 54.8 28.7 23.1 7.0 36.3 16.0
GeoChat [28] 55.1 31.2 28.5 10.1 39.6 19.1
Mini-Gemini [47] 45.2 19.4 17.4 4.6 29.0 10.8
Table 4: Visual grounding performance on VRSBench dataset. Boldface and underline indicate the best and second-best performance.
@@ 4.4 Visual question answering
Evaluation metrics. For the evaluation of our model, we categorize the questions in the test set into 10 distinct types: object category, presence, quantity, color, shape, size, position, direction, scene characteristic, and reasoning. The first eight categories relate to object-level questions, whereas the last two are aligned with scene-level, and reasoning-level questions, respectively. We computed the accuracy for each question type and also determined the average accuracy across all categories.
Results. Table 5 shows the VQA performance of different methods on our VRSBench dataset. As shown in the table, the baseline GeoChat [28] model without finetuning gets an average accuracy of 40.8%. Further finetuning on our VRSBench training set significantly boost the average accuracy to 60.6%. Notably, the visual question-answering capabilities of GPT-4V align closely with those of the optimally tuned GeoChat model, suggesting that this task relies less on specific object information. Qualitative results can be found in the supplementary.
Method Category Presence Quantity Color Shape Size Position Direction Scene Reasoning Avg.
@@
MiniGPT-v2 [38] 46.2 74.1 47.3 44.4 28.6 17.2 23.3 15.3 38.7 36.3 37.1
LLaVA-1.5 [37] 62.8 89.2 50.4 57.8 58.5 52.3 56.9 50.7 66.0 64.9 60.9
GeoChat [28] 60.4 89.9 47.5 58.7 59.1 52.3 57.0 50.3 66.1 64.9 60.6
Mini-Gemini [47] 58.7 89.4 50.0 57.9 57.9 53.7 54.8 50.1 65.0 64.3 60.2
Table 5: Visual question answering performance on VRSBench dataset. Boldface and underline indicate the best and second-best performance.
@@ 7 Broader Impact
By addressing the limitations of existing vision-language datasets, VRSBench provides a comprehensive benchmark for developing and evaluating generalist vision-language models in both remote sensing and computer vision. This dataset not only supports the training and evaluation of advanced vision-language models but also boosts their ability to tackle complex real-world scenarios in remote sensing.
@@ A.2 Data Organizing
  • Images_val.zip contains all raw images in the training split.
  • Annotation_val.zip contains all annotations in the training split, one JSON file per image.
@@ Appendix B Dataset Collection Details
  • GPT-4V annotation generation: The code for GPT-4V prompting is provided at https://github.com/lx709/VRSBench. Detailed instructions are provided in Section F.1.
@@ Appendix E Hosting and Accessibility
The VRSBench dataset is hosted on GitHub (https://github.com/lx709/VRSBench) and Huggingface (https://huggingface.co/datasets/xiang709/VRSBench) to ensure reliable and continuous accessibility.
@@ G.1 Training details
In our experimental setup, all comparative methods are trained on a single node equipped with 4 Nvidia 100 GPUs. The batch size is standardized at 32, and each model undergoes training for a duration of five epochs. We initialize the learning rate at 2e-4 and employ a cosine learning rate decay schedule for optimization. The learning rate experiences a warm-up phase, reaching 3% of the total training steps to gradually adapt to the training regime.
@@ G.2 Visual grounding using OBBs
Settings. In the main paper, horizontal bounding boxes are utilized for both training the model and evaluating its visual grounding capabilities. This section extends the evaluation to incorporate oriented bounding boxes for object localization. Given that GeoChat has demonstrated superior performance in object grounding using bounding boxes, this experiment is exclusively dedicated to exploring the effectiveness of GeoChat under the conditions of oriented bounding boxes. Two distinct configurations of oriented bounding boxes are examined. ‘OBB_1’ is defined by the parameters $[cx,cy,w,h,\theta]$, where $(cx,cy)$ represents the center coordinates, $w$ and $h$ represent the width and height of the bounding box, respectively, and $\theta$ indicates the rotation angle. Conversely, ‘OBB_2’ employs $[x1,y1,x2,y2,\theta]$ for its representation, where $(x1,y1)$ and $(x2,y2)$ denote the coordinates of the top-left and bottom-right points, respectively, with $\theta$ again representing the rotation angle.
Results. From Table 7, the use of $[x1,y1,x2,y2,\theta]$ for representing rotated bounding boxes yields superior performance compared to the $[cx,cy,w,h,\theta]$ representation. Furthermore, the visual grounding performance achieved with oriented bounding boxes is comparable to that observed with horizontal bounding boxes.
@@
Acc@0.5 Acc@0.7 Acc@0.5 Acc@0.7 Acc@0.5 Acc@0.7
GeoChat [28] (OBB_1) 31.7 8.0 16.1 2.9 30.5 5.0
GeoChat [28] (OBB_2) 46.1 13.5 25.5 6.9 34.1 9.6
Table 7: Visual grounding performance on VRSBench dataset using orientated bounding boxes for referring object localization.

### v2

@@ 1 Introduction
Remote sensing models seek to understand the Earth’s surface using imagery captured from overhead, offering a unique perspective of our physical world. This technique is instrumental in various applications, such as land use mapping, urban planning, precision agriculture, disaster management, etc. In the past few years, the success of large vision-language models (LVLMs)[1, 2, 3, 4, 5] in natural scenes has inspired a trend of applying LVLMs to remote sensing [6]. Recent efforts have explored LVLMs for various remote sensing image understanding tasks, including scene classification [7, 8], image captioning [9, 10, 11, 12, 13, 14, 15, 16, 17], visual grounding [18, 19, 20], visual question answering (VQA) [21, 22, 23, 24, 25, 26], and general-purpose models [27, 28], etc.
However, directly applying LVLMs to remote sensing images presents challenges. LVLMs are typically trained on internet data, which differs significantly from remote sensing data. Remote sensing images often feature very small objects (sometimes only 10 pixels) and require complex spatial reasoning from an overhead view. Building effective LVLMs for remote sensing requires large-scale, high-quality datasets tailored to this field. Recent works[27, 28, 29] have attempted to train LVLMs with a combination of existing text-enriched remote sensing data, achieving reasonable performance. However, further improvements are limited by the current vision-language datasets in remote sensing, which have the following limitations:
@@
The key contributions of our work are summarized as follows:
@@
  • VRSBench provides large-scale human-verified annotations that feature several advantages: 1) it provides a large-scale collection of human-verified, high-quality captions rich in object details; 2) it offers more realistic object refers in which each referring sentence unambiguously identifies an object among multiple similar ones within the same category; 3) it features a diverse collection of open-ended question-answer pairs in natural language.
  • We develop three benchmarks based on our VRSBench dataset, including detailed image caption, visual grounding, and visual question answering, and evaluate the performance of several state-of-the-art LVLMs.
@@
#Captions Details #Refers OBB #VQAs Open-ended
@@
Table 1: Comparison between existing remote sensing vision-language datasets and our VRSBench dataset. Values in parentheses in the Caption column indicate the average number of words in captions. OBB denotes orientated bounding box. A small portion of question-answer pairs in RSIVQA are annotated by human annotators.
@@ 2.1 Attribute Extraction
In this study, we utilize two prominent open-access object detection datasets, DOTA-v2 [33] and DIOR [34], to develop our VRSBench dataset. Due to the unavailability of test labels for DOTA-v2, we incorporate only its training and validation sets. We divide each image into patches measuring 512 × 512 pixels. Notably, each image patch from DOTA-v2 contains, on average, 14.2 instances, while each patch from the DIOR dataset averages only 3.3 instances. This higher instance density in DOTA-v2 offers a more challenging and diverse training environment compared to existing remote sensing visual grounding datasets, such as DIOR-RSVG [19] and RRSIS-D [35], that are typically sourced from the DIOR dataset. Moreover, the visual grounding task predominantly involves identifying horizontal bounding boxes (HBB) based on referential descriptions. By constructing our dataset upon the framework of DOTA-v2, VRSBench facilitates the grounding of objects with orientated bounding boxes (OBB), thereby extending the capabilities of traditional visual grounding methods.
@@ 4.1 Benchmark Overview
For the above three tasks, we benchmark state-of-the-art models, including LLaVA-1.5 [37], MiniGPT-v2 [38], and GeoChat [28], to demonstrate the potential of LVLMs for remote sensing image understanding. LLaVA-1.5 [37], MiniGPT-v2 [38], Mini-Gemini [2], and GeoChat [28] are generalist models that are naturally designed for general-proposed image understanding. We, therefore, report the performance of these methods under joint training of all three tasks, i.e., image captioning, visual grounding, and VQA.
Note that all these comparison methods include a multi-stage training process. To ensure a fair comparison, we reload the models that are initially trained on large-scale image-text alignment datasets, and then finetune each method using the training set of our VRSBench dataset. We employ CLIP-ViT(L-14) [39] as the vision encoder and use the Vicuna-7B model [40] as the Large Language Model (LLM). For LLaVA-1.5 [37], Mini-Gemini [2], and GeoChat [28], we adhere to the original model specifications, utilizing two MLP layers with GeLU activation [41]. For MiniGPT-v2[38], we implement a single MLP layer as described in the original paper. For each comparing method, we finetune the model on the training set of VRSBench dataset for 5 epochs. Following GeoChat [28], we use LoRA [42] finetuning to finetune all comparing methods, with a rank of 64. To understand the benefit of fientuning on VRSBench, we include the baseline GeoChat [28] without training on our VRSBench dataset for comparison.
@@ 4.2 Detailed Image Caption
Evaluation metrics. For model evaluation, we follow standard practices by utilizing a set of established metrics including BLEU [43], ROUGE_L [44], METEOR [45], and CIDEr [46]. For BLEU, we consider n-gram precision with n values of 1, 2, 3, and 4. We also report average caption lengths to assess the details of generated captions. Furthermore, we note that traditional caption evaluation metrics may not be suitable for long captions. We, therefore, use an LLM-based caption evaluation metric called CHAIR [^2] https://github.com/davidmchan/clair proposed in [47] for our detailed image caption task.
Results. Table 3 shows the comparative performance of different methods in detailed image captioning of our VRSBench dataset. As demonstrated in the table, the baseline GeoChat model, when not finetuned on the VRSBench dataset, exhibits significantly poorer performance compared to models that have been finetuned on VRSBench. The LLaVA-1.5 [37] model that has undergone fine-tuning on VRSBench achieves the highest performance, reaching a BLEU-1 score of 48.1 and a CIDEr score of 33.9. The GPT-4V model shows the best performance on the CHAIR score. This is expected because the CHIAR score itself is calculated using GPT-4 in our experiments. Moreover, the generated captions have an average word length of 49, which closely approximates the average length of ground truth captions. Note that detailed image captioning is a more challenging task than conventional image caption, therefore, the performance falls far behind. More advanced vision-language modeling techniques are desired to handle this challenging task.
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE_L CIDEr CHAIR Avg_L
GeoChat w/o ft [28] 13.9 6.6 3.0 1.4 7.8 13.2 0.4 0.42 36
GPT-4V [32] 37.2 22.5 13.7 8.6 20.9 30.1 19.1 0.83 67
MiniGPT-v2 [38] 36.8 22.4 13.9 8.7 17.1 30.8 21.4 0.73 37
LLaVA-1.5 [37] 48.1 31.5 21.2 14.7 21.9 36.9 33.9 0.78 49
GeoChat [28] 46.7 30.2 20.1 13.8 21.1 35.2 28.2 0.77 52
Mini-Gemini [48] 47.6 31.1 20.9 14.3 21.5 36.8 33.5 0.77 47
Table 3: Detailed image caption performance on VRSBench dataset. Avg_L denotes the average word length of generated captions. Boldface indicates the best performance.
@@ 4.3 Visual Grounding
Results. Table 3 shows the visual grounding performance of different methods on our VRSBench dataset. From the table, the model finetuned on the VRSBench significantly outperforms the baseline GeoChat model without finetuning. Furthermore, all models demonstrate superior performance in tasks involving unique object referring compared to non-unique object referring. This superiority is expected, as it is generally easier to localize objects uniquely identified within their categories than to differentiate among multiple instances within the same category. Note that even though MiniGPT-v2 gets worse overall grounding performance, it performs better at grounding non-unique objects.
@@
More importantly, even the best-performing GeoChat model fails to achieve satisfactory performance levels, with a grounding accuracy of 49.8% at a threshold of 0.5. This shortfall is attributed to the demanding scenarios presented in the VRSBench dataset, which includes multiple instances of the same category as the target object. This highlights the necessity for more advanced vision grounding techniques to effectively tackle these complexities.
@@
LLaVA-1.5 [37] 51.1 16.4 34.8 11.5 41.6 13.6
GeoChat [28] 57.4 22.6 44.5 18.0 49.8 19.9
Mini-Gemini [48] 41.1 9.6 22.3 4.9 30.1 6.8
Table 4: Visual grounding performance on VRSBench dataset. Boldface indicates the best performance.
@@ 4.4 Visual question answering
Evaluation metrics. We categorize the questions in the test set into 10 distinct types: object category, presence, quantity, color, shape, size, position, direction, scene characteristic, and reasoning. The first eight categories relate to object-level questions, whereas the last two are aligned with scene-level, and reasoning-level questions, respectively. We present the overall accuracy as well as the accuracy for each individual question type. To ensure a robust evaluation, we use GPT-4 to determine for each question whether the answers match ground truth texts, with the prompt: “Question: {question}, Ground Truth Answer: {ground_truth}, Predicted Answer: {predicted answer}. Does the predicted answer match the ground truth? Answer 1 for match and 0 for not match. Use semantic meaning not exact match. Synonyms are also treated as a match, e.g., pond and swimming pool.”.
Results. Table 5 shows the VQA performance of different methods on our VRSBench dataset. As shown in the table, the baseline GeoChat [28] model without finetuning gets an average accuracy of 40.8%. Further finetuning on our VRSBench training set significantly boosts the average accuracy to 76.0%. GPT-4V gets a reasonable performance but still falls a lot behind fine-tuned models, suggesting that detailed object information contributes a lot to the visual question answering task on our benchmark.
Method Category Presence Quantity Color Shape Size Position Direction Scene Reasoning All
@@
MiniGPT-v2 [38] 61.3 26.0 46.1 51.0 41.8 11.2 17.1 12.4 49.3 21.9 38.2
LLaVA-1.5 [37] 86.9 91.8 58.2 69.9 72.2 61.5 69.5 56.7 83.9 73.4 76.4
GeoChat [28] 86.5 92.1 56.3 70.1 73.8 60.4 69.3 53.5 83.7 73.5 76.0
Mini-Gemini [48] 87.8 92.1 58.8 74.0 75.3 58.0 68.0 56.7 83.2 74.4 77.8
Table 5: Visual question answering performance on VRSBench dataset. Boldface indicates the best performance.
@@ 7 Broader Impact
VRSBench provides a comprehensive benchmark for developing and evaluating generalist vision-language models in both remote sensing and computer vision. This dataset not only supports the training and evaluation of advanced vision-language models but also boosts their ability to tackle complex real-world scenarios in remote sensing.
@@ A.2 Data Organizing
  • Images_val.zip contains all raw images in the validation split.
  • Annotation_val.zip contains all annotations in the validation split, one JSON file per image.
@@ Appendix B Dataset Collection Details
  • GPT-4V annotation generation: The code for GPT-4V inference is provided at https://github.com/lx709/VRSBench. Detailed instructions are provided in Section F.1.
@@ Appendix E Hosting and Accessibility
The VRSBench dataset is hosted on Huggingface (https://huggingface.co/datasets/xiang709/VRSBench) to ensure reliable and continuous accessibility.
@@ G.1 Training details
In our experimental setup, all comparative methods are trained on a single node equipped with 4 Nvidia A100 GPUs. The batch size is standardized at 32, and each model undergoes training for a duration of five epochs. We initialize the learning rate at 2e-4 and employ a cosine learning rate decay schedule for optimization. The learning rate experiences a warm-up phase, reaching 3% of the total training steps to gradually adapt to the training regime.
@@ G.2 Visual grounding using OBBs
Settings. In the main paper, horizontal bounding boxes are utilized for both training the model and evaluating its visual grounding capabilities. This section extends the evaluation to incorporate oriented bounding boxes for object localization. Given that GeoChat has demonstrated superior performance in object grounding using bounding boxes, this experiment is exclusively dedicated to exploring the effectiveness of GeoChat under the conditions of oriented bounding boxes. In our experiments, the oriented bounding box is defined by the parameters $[cx,cy,w,h,\theta]$, where $(cx,cy)$ represents the center coordinates, $w$ and $h$ represent the width and height of the bounding box, respectively, and $\theta$ indicates the rotation angle.
Results. As shown in Table 7, the GeoChat model achieves an overall grounding accuracy of 24.3% at a threshold of 0.5, which is lower than its counterpart using horizontal bounding boxes, where the grounding accuracy reaches 49.8%. This highlights that visual grounding with oriented bounding boxes presents a greater challenge compared to grounding with horizontal bounding boxes.
@@
Acc@0.5 Acc@0.7 Acc@0.5 Acc@0.7 Acc@0.5 Acc@0.7 Err_R
GeoChat [28] 32.3 12.6 18.5 5.7 24.3 8.6 18.7
Table 7: Visual grounding performance on VRSBench dataset using orientated bounding boxes for referring object localization. Err_R denotes average angle prediction error.